# MeLA: A Metacognitive LLM-Driven Architecture for Automatic Heuristic Design

### v3

@@ Introduction
To break this impasse, the field of Automatic Heuristic Design (AHD) emerged, aiming to automate the creation of heuristics. While foundational AHD paradigms like Hyper-Heuristics ($H$s) (Pillay and Qu 2018) and Genetic Programming (GP) (Burke et al. 2007) were a conceptual leap forward, their creative potential is fundamentally capped. By design, they are tethered to human guidance, operating within search spaces defined by expert-supplied components or primitives (Mei et al. 2022). This inherent limitation prevents them from achieving true design autonomy, leaving the grand challenge of fully automated discovery unsolved.
@@
The optimization loop begins with an initial problem Analyze, which informs the InitializeHeuristics step, where the first population of heuristics $H$ and their corresponding thought processes ThTh are generated. Subsequently, the core iterative cycle commences. First, all heuristics in the current population are subjected to Evaluation, which yields their fitness scores $f(H)$ and records any execution errors $E$. If $E$ are detected, the CorrectErrors procedure is invoked to repair the faulty heuristics and update the population $H$, thought processes ThTh and fitness $f(H)$. Following this, the AnalyzePerformance stage initiates a metacognitive review MetaMeta. It assesses the recorded thought processes ThTh, the fitness landscape of the population $f(H)$, the errors $E$ and the characteristics of the best-performing heuristic $Best(H)$ to generate a new set of refined prompts. These prompts are then used by the GenerateNext function to produce the next generation of heuristics $H$ and thoughts ThTh. This cycle of evaluation, correction, analysis, and generation repeats, driving the continuous refinement of the underlying problem-solving strategies.
@@
  • Generation Prompts: These prompts create the heuristics. The Initial Generation prompt uses the problem analysis to produce a diverse population of NN heuristics. Subsequently, the Metacognitive Generation prompt uses insights from the metacognition step to evolve this population, generating NN refined heuristics for the next iteration.
@@
Table 1: Comparative Performance Aanalysis Across Different Problems

### v4

@@ Introduction
To break this impasse, the field of Automatic Heuristic Design (AHD) emerged, aiming to automate the creation of heuristics. While foundational AHD paradigms like Hyper-Heuristics (HHs) (Pillay and Qu 2018) and Genetic Programming (GP) (Burke et al. 2007) were a conceptual leap forward, their creative potential is fundamentally capped. By design, they are tethered to human guidance, operating within search spaces defined by expert-supplied components or primitives (Mei et al. 2022). This inherent limitation prevents them from achieving true design autonomy, leaving the grand challenge of fully automated discovery unsolved.
@@
The optimization loop begins with an initial problem Analyze, which informs the InitializeHeuristics step, where the first population of heuristics $H$ and their corresponding thought processes $Th$ are generated. Subsequently, the core iterative cycle commences. First, all heuristics in the current population are subjected to Evaluation, which yields their fitness scores $f(H)$ and records any execution errors $E$. If $E$ are detected, the CorrectErrors procedure is invoked to repair the faulty heuristics and update the population $H$, thought processes $Th$ and fitness $f(H)$. Following this, the AnalyzePerformance stage initiates a metacognitive review $Meta$. It assesses the recorded thought processes $Th$, the fitness landscape of the population $f(H)$, the errors $E$ and the characteristics of the best-performing heuristic $Best(H)$ to generate a new set of refined prompts. These prompts are then used by the GenerateNext function to produce the next generation of heuristics $H$ and thoughts $Th$. This cycle of evaluation, correction, analysis, and generation repeats, driving the continuous refinement of the underlying problem-solving strategies.
@@
  • Generation Prompts: These prompts create the heuristics. The Initial Generation prompt uses the problem analysis to produce a diverse population of $N$ heuristics. Subsequently, the Metacognitive Generation prompt uses insights from the metacognition step to evolve this population, generating $N$ refined heuristics for the next iteration.
@@
Table 1: Comparative Performance Analysis Across Different Problems
@@
In TSP and BPP problems, we optimize the heuristic information based on Ant Colony Optimization (ACO). The core mechanism of ACO involves guiding random solution sampling through heuristic measures. This mechanism is naturally suitable for validating the value of AHD, which automatically designs these heuristic rules to guide the search process, replacing manually designed heuristic functions(Ye et al. 2024). In ACS and WSN problems, we optimize the heuristic algorithm. By evaluating the generated results of heuristic information and heuristic algorithms, we assess the excellent performance of MeLA.
@@
It should be emphasized that for excellent algorithms such as GA, in the TSP and BPP problems, we evaluate the ACO’s performance with heuristic information designed by these algorithms, while in the ACS and BPP problems, we evaluate the algorithms themselves.